OpenStudy (anonymous):

use the method of substitution to solve the system of linear equations 3x-y=2 7x-2y=8

OpenStudy (mathstudent55):

Solve the first equation for y. Then substitute in second equation.

OpenStudy (anonymous):

im still confused

OpenStudy (anonymous):

so the second answer would be done the same

OpenStudy (mathstudent55):

The first equation is: 3x - y = 2 Since y is not being multiplied by a number, it's easy to solve that equation for y: 3x = y + 2 3x - 2 = y Now that we know y = 3x - 2, we substitute that for y in the second equation: 7x - 2y = 8 7x -2(3x - 2) = 8 7x - 6x + 4 = 8 x + 4 = 8 x = 4 Now we substitute the value of x into the first equation to find y: 3x - y = 2 3(4) - y = 2 12 - y = 2 -y = -10 y = 10 The solution is x = 4 and y = 10

OpenStudy (mathstudent55):

@Kenneditodd wrote: "3x-y=2 -y=-3x+2 y=3x-2 then substitute y into the second equation 7x-2y=8 x-2(3x-2)=8 x-6x+4=8 -5x+4=8 -5x=4 x=-4/5" From the line 7x-2y=8 to the line x-2(3x-2)=8 you brought down x instead of 7x. That's why your answer is incorrect.

OpenStudy (anonymous):

thanks! oops!

OpenStudy (mathstudent55):

No problem. It happens to all of us.

OpenStudy (anonymous):

so what is the correct answer may I ask

OpenStudy (mathstudent55):

4, 10

OpenStudy (anonymous):

thanks

OpenStudy (mathstudent55):

wlcm